the function f(x) = modulus of (a*x-b) + c*modulus of (x) for all x belongs to real number, where a,b,c > 0 assumes it's minimum only at one point. Find that condition.

Grade:12th pass

1 Answers

Jatin Garg
6 years ago
the function f(x) = modulus of (a*x-b) + c*modulus of (x) for all x belongs to real number, where a,b,c > 0 will give one minimum value if and only if c is not equal to aFor x b/a, f(x)=(a+c)x -b, Inc function for all x with min value =(b×c)/a,Hence, for the function to give just one minimum value c should not be equal to a, if it`s so the function will give a min value for a range of values.